# Character Theory for Semilinear Representations

### v1

@@ Abstract.
Let $G$ be a group acting on a field $L$, and suppose that $L/K$ is a finite Galois extension, where $K=L^{G}$. We show that the irreducible semilinear representations of $G$ over $L$ can be completely described in terms of irreducible linear representations of $H$, the kernel of $G\rightarrow\operatorname{Gal}(L/K)$. When $G$ is finite and $|G|\in L^{\times}$ this provides a character theory for semilinear representations of $G$ over $L$, which recovers ordinary character theory when the action of $G$ on $L$ is trivial.
@@ 1.2. The Extension Problem
In particular, the statement that there is at most one extension of $W$ to a semilinear representation of $C_{4}$ over ${\mathbb{Q}}(\sqrt{d})$ if equivalent to the claim that if $X^{2}-dY^{2}=-1$ has a solution $(x,y)\in{\mathbb{Q}}^{2}$, then the set of all solutions in ${\mathbb{Q}}^{2}$ is parametrised by
$\left(\frac{(u^{2}+dv^{2})x+(2uv)dy}{u^{2}-dv^{2}},\frac{(u^{2}+dv^{2})y+(2uv)x}{u^{2}-dv^{2}}\right)$
over $u:v]\in{\mathbb{P}}^{1}({\mathbb{Q}})$.
@@ 1.3. Contribution
In this paper we provide what we believe to be the first systematic description of the semilinear representation categories $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ beyond the linear case of $\operatorname{\mathbf{Rep}}_{L}(G)$. We work under the following assumption, which we impose from now until the end of the introduction:
  • $L/K$ is a finite Galois extension.
@@ 1.5. Group Action
This defines an left action of $G$ on $\operatorname{\mathbf{Rep}}_{L}(H)/\!\cong$ which is trivial on $H$, and therefore induces an action of the quotient $G/H$. The induced map $G\rightarrow\operatorname{Gal}(L/K)$ is automatically surjective (Remark 3.5), and therefore the this induces an action of $\operatorname{Gal}(L/K)$ on $\operatorname{\mathbf{Rep}}_{L}(H)/\!\cong$.
If $W=V|_{H}$ for some $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, then for any $g\in G$, the action of $g$ on $V$ defines an isomorphism $\rho(g)\colon g*W\xrightarrow{\sim}W$, and so $W$ is fixed by $\operatorname{Gal}(L/K)$. So a necessary condition for $W$ to extend to a semilinear representation of $G$ over $L$ is that $W$ is fixed by $\operatorname{Gal}(L/K)$.
@@ 1.6. Irreducible Representations
we also consider the induction functor
@@ 1.10. Determination of Semilinear Schur Indices
For classical Schur indices, one has Ungers algorithm [UNG] for computing Schur indices over number fields, and it would be interesting to see if this can be generalised to semilinear Schur indices. There is still a local-global principal here (Corollary 11.5), and so the problem can similarly be reduced to a problem for local fields. However many results upon which the algorithm rests require deep facts about classical Schur indices, which at present are not established for semilinear Schur indices.
@@ 1.11. Structure of the Paper
In Section 11 we study the splitting behaviour of the $K$-algebras $\operatorname{End}_{L\rtimes G}(V)$ for $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and consider base change for semilinear representations. When $K$ is a number field we use the local-global principal for $\operatorname{Br}(K)$ to prove a local-global principal for semilinear representations.
@@ Acknowledgements
I would like to thank Tom Adams, Lorenzo La Porta and Dmitriy Rumynin for interesting conversations regarding the contents of this paper. This research was supported by an LMS Early Career Research Fellowship at the University of Cambridge and the departmental grant Progetto Sviluppo Dipartimentale - UNIPD PSDIP23O88 at the University of Padova.
@@ 3.1. Galois Extensions of a Field
Generally, if $L/K$ is a Galois extension with Galois group $\Gamma$, then for any connected component $\operatorname{Spec}(L_{i})$ of $\operatorname{Spec}(L)$, $L_{i}/K$ is a Galois field extension with Galois group $G_{i}\coloneqq\operatorname{Stab}_{\Gamma}(\operatorname{Spec}(L_{i}))$, and the action of $\Gamma$ on $\pi_{0}(\operatorname{Spec}(L))$ is transitive [TAY3, Lem. 2.51].
@@ 3.2. Assumptions and Notation
Suppose momentarily that $L$ is a field and $G$ is a group that acts on $L$ through a group homomorphism $G\rightarrow\operatorname{Aut}(L)$ (as in the introduction). Setting $K\coloneqq L^{G}$, then the assumption that $L/K$ is a finite Galois extension automatically implies that the induced map $G\rightarrow\operatorname{Gal}(L/K)$ is surjective (and therefore satisfies the assumptions of this section). Indeed, if $S\subset\operatorname{Gal}(L/K)$ denotes the image of $G$, then $L^{S}=L^{G}=K$, and so $S=\operatorname{Gal}(L/K)$ by the Galois correspondence.
@@ 6. Split Extensions
This is just Galois descent, see for example [TAY3, Prop. 2.52]. ∎
@@ 10.1.2. Divisibility by Dimension
The following result implies that any linear character of $H$ over $L$ has Schur index $1$. This makes sense, as any such character fixed by $\operatorname{Gal}(L/K)$ is defined over $K$.
@@
Suppose that $H$ is finite, and $|H|\in L^{\times}$. Suppose that $W\in\operatorname{Irr}_{L}(H)$, and let $D_{W}\coloneqq\operatorname{End}_{L[H]}(W)$, a (potentially non-central) division algebra over $L$. Then
@@
Under these assumptions, $K[H]$ is semisimple, and therefore by Corollary 10.3(4), $\overline{m}_{K}^{L}(W)$ divides the multiplicity of $W$ in $K[H]_{L}=L[H]$, which is exactly $\dim_{L}(W)/\dim_{L}D_{W}$. ∎
The same proof does not work in the non-split setting, if one tries to use the algebra $L\rtimes G$ instead of $K[H]$. Furthermore, the statement for semilinear Schur indices is false, as Example 9.5 shows, taking $L/K$ to be ${\mathbb{Q}}(i)/{\mathbb{Q}}$.
@@ 11. Base Change for Semilinear Representations
We can use this to give a local-global principal in our context.

### v2

@@ Abstract.
Let $G$ be a group acting on a field $L$, and suppose that $L/L^{G}$ is a finite extension. We show that the irreducible semilinear representations of $G$ over $L$ can be completely described in terms of irreducible linear representations of $H$, the kernel of the map $G\rightarrow\operatorname{Aut}(L)$. When $G$ is finite and $|G|\in L^{\times}$ this provides a character theory for semilinear representations of $G$ over $L$, which recovers ordinary character theory when the action of $G$ on $L$ is trivial.
@@ 1.2. The Extension Problem
In particular, the statement that there is at most one extension of $W$ to a semilinear representation of $C_{4}$ over ${\mathbb{Q}}(\sqrt{d})$ if equivalent to the claim that if $X^{2}-dY^{2}=-1$ has a solution $(x,y)\in{\mathbb{Q}}^{2}$, then the set of all solutions in ${\mathbb{Q}}^{2}$ is parametrised by $u:v]\in{\mathbb{P}}^{1}({\mathbb{Q}})$ via the formula
$\left(\frac{(u^{2}+dv^{2})x+(2uv)dy}{u^{2}-dv^{2}},\frac{(u^{2}+dv^{2})y+(2uv)x}{u^{2}-dv^{2}}\right).$
@@ 1.3. Contribution
In this paper we provide what we believe to be the first systematic description of the semilinear representation categories $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ beyond the linear case of $\operatorname{\mathbf{Rep}}_{L}(G)$. We work under the following mild assumption, which we impose from now until the end of the introduction:
  • $L/K$ is a finite extension.
For example, this is automatically satisfied if either $G$ is finite or $L$ is finite over its prime subfield. We also note this further implies that $L/K$ is Galois and $G\rightarrow\operatorname{Gal}(L/K)$ is surjective (Remark 3.5).
@@ 1.5. Group Action
This defines an left action of $G$ on $\operatorname{\mathbf{Rep}}_{L}(H)/\!\cong$ which is trivial on $H$, and therefore induces an action of the quotient $G/H$, and so also $\operatorname{Gal}(L/K)$ using the isomorphism $G/H\xrightarrow{\sim}\operatorname{Gal}(L/K)$.
If $W=V|_{H}$ for some $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, then for any $g\in G$, the action of $g$ on $V$ defines an isomorphism $\rho(g)\colon g*W\xrightarrow{\sim}W$, and so $W$ is fixed by $\operatorname{Gal}(L/K)$. Therefore a necessary condition for $W$ to extend to a semilinear representation of $G$ over $L$ is that $W$ is fixed by $\operatorname{Gal}(L/K)$.
@@ 1.6. Irreducible Representations
we can, denoting by $L\rtimes G$ the skew group ring (Definition 2.9), consider the induction functor
@@ 1.10. Determination of Semilinear Schur Indices
For classical Schur indices, one has Unger’s algorithm [UNG] for computing Schur indices over number fields, and it would be interesting to see if this can be generalised to semilinear Schur indices. There is still a local-global principle here (Corollary 11.5), and so the problem can similarly be reduced to a problem for local fields. However many results upon which the algorithm rests require deep facts about classical Schur indices, which at present are not established for semilinear Schur indices.
@@ 1.11. Structure of the Paper
In Section 11 we study the splitting behaviour of the $K$-algebras $\operatorname{End}_{L\rtimes G}(V)$ for $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and consider base change for semilinear representations. When $K$ is a number field we use the local-global principle for $\operatorname{Br}(K)$ to prove a local-global principle for semilinear representations.
@@ Acknowledgements
I would like to thank Tom Adams, Håvard Damm-Johnsen, Lorenzo La Porta and Dmitriy Rumynin for interesting conversations regarding the contents of this paper. This research was supported by an LMS Early Career Research Fellowship at the University of Cambridge and the departmental grant Progetto Sviluppo Dipartimentale - UNIPD PSDIP23O88 at the University of Padova.
@@ 3.1. Galois Extensions of a Field
Generally, if $L/K$ is a Galois extension with Galois group $\Gamma$, then for any connected component $\operatorname{Spec}(L_{i})$ of $\operatorname{Spec}(L)$, $L_{i}/K$ is a Galois field extension with Galois group $G_{i}\coloneqq\operatorname{Stab}_{\Gamma}(\operatorname{Spec}(L_{i}))$, and the action of $\Gamma$ on $\pi_{0}(\operatorname{Spec}(L))$ is transitive [TAY3, Lem. 2.52].
@@ 3.2. Assumptions and Notation
Suppose momentarily that $L$ is a field and $G$ is a group that acts on $L$ through a group homomorphism $G\rightarrow\operatorname{Aut}(L)$ (as in the introduction). Setting $K\coloneqq L^{G}$, then the assumption that $L/K$ is a finite extension automatically implies that $L/K$ is Galois and that the induced map $G\rightarrow\operatorname{Gal}(L/K)$ is surjective (and therefore satisfies the assumptions of this section). Indeed, $K$ is the fixed field of a finite subgroup of $\operatorname{Aut}(L)$, so $L/K$ is Galois, and if $S\subset\operatorname{Gal}(L/K)$ denotes the image of $G$ then $L^{S}=L^{G}=K$, hence $S=\operatorname{Gal}(L/K)$ by the Galois correspondence.
@@ 6. Split Extensions
This is just Galois descent, see for example [TAY3, Prop. 2.53]. ∎
@@ 10.1.2. Divisibility by Dimension
The following result, amongst other things, implies that any linear character of $H$ over $L$ has Schur index $1$. This makes sense, as any such representation fixed by $\operatorname{Gal}(L/K)$ is easily seen to be defined over $K$. The proof does not extend to the non-split setting, and indeed the statement for semilinear Schur indices is false. This can be seen from Example 9.5, taking $L/K$ to be ${\mathbb{Q}}(i)/{\mathbb{Q}}$, which exhibits linear characters of $H$ over $L$ with Schur index $2$.
@@
Suppose that $W\in\operatorname{Irr}_{L}(H)$, and let $D_{W}\coloneqq\operatorname{End}_{L[H]}(W)$, a (potentially non-central) division algebra over $L$. Then
@@
Let $V\in\operatorname{Irr}_{K}(H)$ correspond to $W$, and set $m\coloneqq\overline{m}_{K}^{L}$. Writing $D_{V}\coloneqq\operatorname{End}_{K[G]}(V)$, we have that $\dim_{K}D_{V}\mid\dim_{K}V$ as $V$ is a free module over $D_{V}$. From Corollary 10.1 we have that
$V_{L}=\oplus_{\gamma\in\Gamma/\Gamma_{W}}(\gamma*W)^{m}$
and thus
$\dim_{K}D_{V}=\dim_{L}\operatorname{End}_{L[H]}(V_{L})=\dim_{L}\prod_{\Gamma/\Gamma_{W}}M_{m}(D_{W})=[\Gamma:\Gamma_{W}]\cdot m^{2}\cdot\dim_{L}D_{W}.$
From $\dim_{K}D_{V}\mid\dim_{K}V$ and noting that $\dim_{K}V=\dim_{L}V_{L}$ we have that
$\Gamma:\Gamma_{W}]\cdot m^{2}\cdot\dim_{L}D_{W}\mid[\Gamma:\Gamma_{W}]\cdot m\cdot\dim_{L}W,$
and therefore $m\cdot\dim_{L}D_{W}\mid\dim_{L}W$ as required. ∎
@@ 11. Base Change for Semilinear Representations
We can use this to give a local-global principle in our context.